A wheel starts from rest and has an angular acceleration of 4.0 rad/s2. What time does it takes to make 100rev?
1
Expert's answer
2020-06-12T11:17:46-0400
"\theta=0.5\alpha t^2"

"100(2\pi)=0.5(4)t^2"

"t=17.7\ s"